X3J13 Document 88-002R June 1988

Common Lisp Object System Specification

1. Programmer Interface Concepts

Authors: Daniel G. Bobrow, Linda G. DeMichiel, Richard P. Gabriel, Sonya E. Keene, Gregor Kiczales, and David A. Moon.

Draft Dated: June 15, 1988 All Rights Reserved

The distribution and publication of this document are not restricted. In order to preserve the integrity of the specification, any publication or distribution must reproduce this document in its entirety, preserve its formatting, and include this title page.

For information about obtaining the sources for this document, send an Internet message to common-lisp-object-system-specification-request@sail.stanford.edu.

The authors wish to thank Patrick Dussud, Kenneth Kahn, Jim Kempf, Larry Masinter, Mark Stefik, Daniel L. Weinreb, and Jon L White for their contributions to this document.

At the X3J13 meeting on June 15, 1988, the following motion was adopted:

“The X3J13 Committee hereby accepts chapters 1 and 2 of the Common Lisp Object System, as defined in document 88-002R, for inclusion in the Common Lisp language being specified by this committee. Subsequent changes will be handled through the usual editorial and cleanup processes.”

CONTENTS

Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–3 Error Terminology . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–4 Classes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–6

Defining Classes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–7 Creating Instances of Classes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–8 Slots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–8 Accessing Slots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–9

Inheritance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–10 Inheritance of Methods . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–10 Inheritance of Slots and Slot Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–10 Inheritance of Class Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–11 Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–12

Integrating Types and Classes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–13 Determining the Class Precedence List . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–16

Topological Sorting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–16 Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–17

Generic Functions and Methods . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–19 Introduction to Generic Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–19 Introduction to Methods . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–20 Agreement on Parameter Specializers and Qualifiers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–22 Congruent Lambda-Lists for All Methods of a Generic Function . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–23 Keyword Arguments in Generic Functions and Methods . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–23

Method Selection and Combination . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–25

Determining the Effective Method . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–25 Selecting the Applicable Methods . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–25 Sorting the Applicable Methods by Precedence Order . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–25 Applying Method Combination to the Sorted List of Applicable Methods . . . . . . . . . . . . . . . . . . 1–26

Standard Method Combination . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–27 Declarative Method Combination . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–29 Built-in Method Combination Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–29

Meta-Objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–31 Metaclasses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–31 Standard Metaclasses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–31 Standard Meta-objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–31

Object Creation and Initialization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–33 Initialization Arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–34 Declaring the Validity of Initialization Arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–34 Defaulting of Initialization Arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–35 Rules for Initialization Arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–36 Shared-Initialize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–37 Initialize-Instance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–38 Definitions of Make-Instance and Initialize-Instance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–39

Redefining Classes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–41 Modifying the Structure of Instances . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–41 Initializing Newly Added Local Slots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–42 Customizing Class Redefinition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–42 Extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–43

Changing the Class of an Instance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–44 Modifying the Structure of the Instance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–44 Initializing Newly Added Local Slots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–44 Customizing the Change of Class of an Instance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–45

Reinitializing an Instance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–46 Customizing Reinitialization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1–46

Introduction

The Common Lisp Object System is an object-oriented extension to Common Lisp as defined in Common Lisp: The Language, by Guy L. Steele Jr. It is based on generic functions, multiple inheritance, declarative method combination, and a meta-object protocol.

The first two chapters of this specification present a description of the standard Programmer Interface for the Common Lisp Object System. The first chapter contains a description of the concepts of the Common Lisp Object System, and the second contains a description of the functions and macros in the Common Lisp Object System Programmer Interface. The chapter “The Common Lisp Object System Meta-Object Protocol” describes how the Common Lisp Object System can be customized.

The fundamental objects of the Common Lisp Object System are classes, instances, generic functions, and methods.

A class object determines the structure and behavior of a set of other objects, which are called its instances. Every Common Lisp object is an instance of a class. The class of an object determines the set of operations that can be performed on the object.

A generic function is a function whose behavior depends on the classes or identities of the arguments supplied to it. A generic function object contains a set of methods, a lambda-list, a method combination type, and other information. The methods define the class-specific behavior and operations of the generic function; a method is said to specialize a generic function. When invoked, a generic function executes a subset of its methods based on the classes of its arguments.

A generic function can be used in the same ways that an ordinary function can be used in Common Lisp; in particular, a generic function can be used as an argument to funcall and apply and can be given a global or a local name.

A method is an object that contains a method function, a sequence of parameter specializers that specify when the given method is applicable, and a sequence of qualifiers that is used by the method combination facility to distinguish among methods. Each required formal parameter of each method has an associated parameter specializer, and the method will be invoked only on arguments that satisfy its parameter specializers.

The method combination facility controls the selection of methods, the order in which they are run, and the values that are returned by the generic function. The Common Lisp Object System offers a default method combination type and provides a facility for declaring new types of method combination.

Error Terminology

The terminology used in this document to describe erroneous situations differs from the terminology used in Common Lisp: The Language, by Guy L. Steele Jr. This terminology involves situations; a situation is the evaluation of an expression in some specific context. For example, a situation might be the invocation of a function on arguments that fail to satisfy some specified constraints.

In the specification of the Common Lisp Object System, the behavior of programs in all situations is described, and the options available to the implementor are defined. No implementation is allowed to extend the syntax or semantics of the Object System except as explicitly defined in the Object System specification. In particular, no implementation is allowed to extend the syntax of the Object System in such a way that ambiguity between the specified syntax of Object System and those extensions is possible.

“When situation S occurs, an error is signaled.”

This terminology has the following meaning:

• If this situation occurs, an error will be signaled in the interpreter and in code compiled under all compiler safety optimization levels.

• Valid programs may rely on the fact that an error will be signaled in the interpreter and in code compiled under all compiler safety optimization levels.

• Every implementation is required to detect such an error in the interpreter and in code compiled under all compiler safety optimization levels.

“When situation S occurs, an error should be signaled.”

This terminology has the following meaning:

• If this situation occurs, an error will be signaled at least in the interpreter and in code compiled under the safest compiler safety optimization level.

• Valid programs may not rely on the fact that an error will be signaled.

• Every implementation is required to detect such an error at least in the interpreter and in code compiled under the safest compiler safety optimization level.

• When an error is not signaled, the results are undefined (see below).

“When situation S occurs, the results are undefined.”

This terminology has the following meaning:

• If this situation occurs, the results are unpredictable. The results may range from harmless to fatal.

• Implementations are allowed to detect this situation and signal an error, but no implementation is required to detect the situation.

• No valid program may depend on the effects of this situation, and all valid programs are required to treat the effects of this situation as unpredictable.

“When situation S occurs, the results are unspecified.” This terminology has the following meaning:

• The effects of this situation are not specified in the Object System, but the effects are harmless.

• Implementations are allowed to specify the effects of this situation.

• No portable program can depend on the effects of this situation, and all portable programs are required to treat the situation as unpredictable but harmless.

“The Common Lisp Object System may be extended to cover situation S.” The meaning of this terminology is that an implementation is free to treat situation S in one of three ways:

• When situation S occurs, an error is signaled at least in the interpreter and in code compiled under the safest compiler safety optimization level.

• When situation S occurs, the results are undefined.

• When situation S occurs, the results are defined and specified. In addition, this terminology has the following meaning:

• No portable program can depend on the effects of this situation, and all portable programs are required to treat the situation as undefined.

“Implementations are free to extend the syntax S.” This terminology has the following meaning:

• Implementations are allowed to define unambiguous extensions to syntax S.

• No portable program can depend on this extension, all portable programs are required to treat the syntax as meaningless.

The Common Lisp Object System specification may disallow certain extensions while allowing others.

Classes

A class is an object that determines the structure and behavior of a set of other objects, which are called its instances.

A class can inherit structure and behavior from other classes. A class whose definition refers to other classes for the purpose of inheriting from them is said to be a subclass of each of those classes. The classes that are designated for purposes of inheritance are said to be superclasses of the inheriting class.

A class can have a name. The function class-name takes a class object and returns its name. The name of an anonymous class is nil. A symbol can name a class. The function find-class takes a symbol and returns the class that the symbol names. A class has a proper name if the name is a symbol and if the name of the class names that class. That is, a class C has the proper name S if S = (class-name C) and C = (find-class S). Notice that it is possible for

(find-class S_{1}) = (find-class S_{2}) and S_{1} ≠ S_{2}. If C = (find-class S), we say that C is the

class named S.

A class C_{1} is a direct superclass of a class C_{2} if C_{2} explicitly designates C_{1} as a superclass in its definition. In this case C_{2} is a direct subclass of C_{1}. A class C_{n} is a superclass of a class C_{1} if there exists a series of classes C_{2}, …, C_{n−1} such that C_{i+1} is a direct superclass of C_{i} for 1 ≤ i < n. In this case, C_{1} is a subclass of C_{n}. A class is considered neither a superclass nor a subclass of itself. That is, if C_{1} is a superclass of C_{2}, then C_{1} ≠ C_{2}. The set of classes consisting of some given class C along with all of its superclasses is called “C and its superclasses.”

Each class has a class precedence list, which is a total ordering on the set of the given class and its superclasses. The total ordering is expressed as a list ordered from most specific to least specific. The class precedence list is used in several ways. In general, more specific classes can shadow, or override, features that would otherwise be inherited from less specific classes. The method selection and combination process uses the class precedence list to order methods from most specific to least specific.

When a class is defined, the order in which its direct superclasses are mentioned in the defining form is important. Each class has a local precedence order, which is a list consisting of the class followed by its direct superclasses in the order mentioned in the defining form.

A class precedence list is always consistent with the local precedence order of each class in the list. The classes in each local precedence order appear within the class precedence list in the same order. If the local precedence orders are inconsistent with each other, no class precedence list can be constructed, and an error is signaled. The class precedence list and its computation is discussed in the section “Determining the Class Precedence List.”

Classes are organized into a directed acyclic graph. There are two distinguished classes, named t and standard-object. The class named t has no superclasses. It is a superclass of every class except itself. The class named standard-object is an instance of the class standard-class and is a superclass of every class that is an instance of standard-class except itself.

There is a mapping from the Common Lisp Object System class space into the Common Lisp type space. Many of the standard Common Lisp types specified in Common Lisp: The Language have a corresponding class that has the same name as the type. Some Common Lisp types do not have a corresponding class. The integration of the type and class systems is discussed in the section “Integrating Types and Classes.”

Classes are represented by objects that are themselves instances of classes. The class of the class of an object is termed the metaclass of that object. When no misinterpretation is possible, the term metaclass will be used to refer to a class that has instances that are themselves classes. The metaclass determines the form of inheritance used by the classes that are its instances and the representation of the instances of those classes. The Common Lisp Object System provides a default metaclass, standard-class, that is appropriate for most programs. The meta-object protocol provides mechanisms for defining and using new metaclasses.

Except where otherwise specified, all classes mentioned in this chapter are instances of the class standard-class, all generic functions are instances of the class standard-generic-function, and all methods are instances of the class standard-method.

Defining Classes

The macro defclass is used to define a new named class. The syntax for defclass is given in Figure 2-1.

The definition of a class includes:

• The name of the new class. For newly defined classes this name is a proper name.

• The list of the direct superclasses of the new class.

• A set of slot specifiers. Each slot specifier includes the name of the slot and zero or more slot options. A slot option pertains only to a single slot. If a class definition contains two slot specifiers with the same name, an error is signaled.

• A set of class options. Each class option pertains to the class as a whole.

The slot options and class options of the defclass form provide mechanisms for the following:

• Supplying a default initial value form for a given slot.

• Requesting that methods for generic functions be automatically generated for reading or writing slots.

• Controlling whether a given slot is shared by instances of the class or whether each instance of the class has its own slot.

• Supplying a set of initialization arguments and initialization argument defaults to be used in instance creation.

• Indicating that the metaclass is to be other than the default.

• Indicating the expected type for the value stored in the slot.

• Indicating the documentation string for the slot.

Creating Instances of Classes

The generic function make-instance creates and returns a new instance of a class. The Object System provides several mechanisms for specifying how a new instance is to be initialized. For example, it is possible to specify the initial values for slots in newly created instances either by giving arguments to make-instance or by providing default initial values. Further initialization activities can be performed by methods written for generic functions that are part of the initialization protocol. The complete initialization protocol is described in the section “Object Creation and Initialization.”

Slots

An object that has standard-class as its metaclass has zero or more named slots. The slots of an object are determined by the class of the object. Each slot can hold one value. The name of a slot is a symbol that is syntactically valid for use as a Common Lisp variable name.

When a slot does not have a value, the slot is said to be unbound. When an unbound slot is read, the generic function slot-unbound is invoked. The system-supplied primary method for slot-unbound signals an error.

The default initial value form for a slot is defined by the :initform slot option. When the :initform form is used to supply a value, it is evaluated in the lexical environment in which the defclass form was evaluated. The :initform along with the lexical environment in which the defclass form was evaluated is called a captured :initform. See the section “Object Creation and Initialization” for more details.

A local slot is defined to be a slot that is visible to exactly one instance, namely the one in which the slot is allocated. A shared slot is defined to be a slot that is visible to more than one instance of a given class and its subclasses.

A class is said to define a slot with a given name when the defclass form for that class contains a slot specifier with that name. Defining a local slot does not immediately create a slot; it causes a slot to be created each time an instance of the class is created. Defining a shared slot immediately creates a slot.

The :allocation slot option to defclass controls the kind of slot that is defined. If the value of the :allocation slot option is :instance, a local slot is created. If the value of :allocation is :class, a shared slot is created.

A slot is said to be accessible in an instance of a class if the slot is defined by the class of the instance or is inherited from a superclass of that class. At most one slot of a given name can be accessible in an instance. A shared slot defined by a class is accessible in all instances of that class. A detailed explanation of the inheritance of slots is given in the section “Inheritance of Slots and Slot Options.”

Accessing Slots

Slots can be accessed in two ways: by use of the primitive function slot-value and by use of generic functions generated by the defclass form.

The function slot-value can be used with any of the slot names specified in the defclass form to access a specific slot accessible in an instance of the given class.

The macro defclass provides syntax for generating methods to read and write slots. If a reader is requested, a method is automatically generated for reading the value of the slot, but no method for storing a value into it is generated. If a writer is requested, a method is automatically generated for storing a value into the slot, but no method for reading its value is generated. If an accessor is requested, a method for reading the value of the slot and a method for storing a value into the slot are automatically generated. Reader and writer methods are implemented using slot-value.

When a reader or writer is specified for a slot, the name of the generic function to which the generated method belongs is directly specified. If the name specified for the writer option is the symbol name, the name of the generic function for writing the slot is the symbol name, and the generic function takes two arguments: the new value and the instance, in that order. If the name specified for the accessor option is the symbol name, the name of the generic function for reading the slot is the symbol name, and the name of the generic function for writing the slot is the list

(setf name).

A generic function created or modified by supplying reader, writer, or accessor slot options can be treated exactly as an ordinary generic function.

Note that slot-value can be used to read or write the value of a slot whether or not reader or writer methods exist for that slot. When slot-value is used, no reader or writer methods are invoked.

The macro with-slots can be used to establish a lexical environment in which specified slots are lexically available as if they were variables. The macro with-slots invokes the function slot-value to access the specified slots.

The macro with-accessors can be used to establish a lexical environment in which specified slots are lexically available through their accessors as if they were variables. The macro withaccessors invokes the appropriate accessors to access the specified slots. Any accessors specified by with-accessors must already have been defined before they are used.

Inheritance

A class can inherit methods, slots, and some defclass options from its superclasses. The following sections describe the inheritance of methods, the inheritance of slots and slot options, and the inheritance of class options.

Inheritance of Methods

A subclass inherits methods in the sense that any method applicable to all instances of a class is also applicable to all instances of any subclass of that class.

The inheritance of methods acts the same way regardless of whether the method was created by using one of the method-defining forms or by using one of the defclass options that causes methods to be generated automatically.

The inheritance of methods is described in detail in the section “Method Selection and Combination.”

Inheritance of Slots and Slot Options

The set of the names of all slots accessible in an instance of a class C is the union of the sets of names of slots defined by C and its superclasses. The structure of an instance is the set of names of local slots in that instance.

In the simplest case, only one class among C and its superclasses defines a slot with a given slot name. If a slot is defined by a superclass of C, the slot is said to be inherited. The characteristics of the slot are determined by the slot specifier of the defining class. Consider the defining class for a slot S. If the value of the :allocation slot option is :instance, then S is a local slot and each instance of C has its own slot named S that stores its own value. If the value of the :allocation slot option is :class, then S is a shared slot, the class that defined S stores the value, and all instances of C can access that single slot. If the :allocation slot option is omitted, :instance is used.

In general, more than one class among C and its superclasses can define a slot with a given name. In such cases, only one slot with the given name is accessible in an instance of C, and the characteristics of that slot are a combination of the several slot specifiers, computed as follows:

• All the slot specifiers for a given slot name are ordered from most specific to least specific, according to the order in C’s class precedence list of the classes that define them. All references to the specificity of slot specifiers immediately below refers to this ordering.

• The allocation of a slot is controlled by the most specific slot specifier. If the most specific slot specifier does not contain an :allocation slot option, :instance is used. Less specific slot specifiers do not affect the allocation.

• The default initial value form for a slot is the value of the :initform slot option in the most specific slot specifier that contains one. If no slot specifier contains an :initform slot option, the slot has no default initial value form.

• The contents of a slot will always be of type (and T_{1} … T_{n}) where T_{1} … T_{n} are the values of the :type slot options contained in all of the slot specifiers. If no slot specifier contains the :type slot option, the contents of the slot will always be of type t. The result of attempting to store in a slot a value that does not satisfy the type of the slot is undefined.

• The set of initialization arguments that initialize a given slot is the union of the initialization arguments declared in the :initarg slot options in all the slot specifiers.

• The documentation string for a slot is the value of the :documentation slot option in the most specific slot specifier that contains one. If no slot specifier contains a :documentation slot option, the slot has no documentation string.

A consequence of the allocation rule is that a shared slot can be shadowed. For example, if a class C_{1} defines a slot named S whose value for the :allocation slot option is :class, that slot is accessible in instances of C_{1} and all of its subclasses. However, if C_{2} is a subclass of C_{1} and also defines a slot named S, C_{1}’s slot is not shared by instances of C_{2} and its subclasses. When a class C_{1} defines a shared slot, any subclass C_{2} of C_{1} will share this single slot unless the defclass form for C_{2} specifies a slot of the same name or there is a superclass of C_{2} that precedes C_{1} in the class precedence list of C_{2} that defines a slot of the same name.

A consequence of the type rule is that the value of a slot satisfies the type constraint of each slot specifier that contributes to that slot. Because the result of attempting to store in a slot a value that does not satisfy the type constraint for the slot is undefined, the value in a slot might fail to satisfy its type constraint.

The :reader, :writer, and :accessor slot options create methods rather than define the characteristics of a slot. Reader and writer methods are inherited in the sense described in the section “Inheritance of Methods.”

Methods that access slots use only the name of the slot and the type of the slot’s value. Suppose a superclass provides a method that expects to access a shared slot of a given name, and a subclass defines a local slot with the same name. If the method provided by the superclass is used on an instance of the subclass, the method accesses the local slot.

Inheritance of Class Options

The :default-initargs class option is inherited. The set of defaulted initialization arguments for a class is the union of the sets of initialization arguments specified in the :default-initargs class options of the class and its superclasses. When more than one default initial value form is supplied for a given initialization argument, the default initial value form that is used is the one supplied by the class that is most specific according to the class precedence list.

If a given :default-initargs class option specifies an initialization argument of the same name more than once, an error is signaled.

Examples

(defclass C1 ()
    ((S1 :initform 5.4 :type number)
     (S2 :allocation :class)))
(defclass C2 (C1)
    ((S1 :initform 5 :type integer)
     (S2 :allocation :instance)
     (S3 :accessor C2-S3)))

Instances of the class C1 have a local slot named S1, whose default initial value is 5.4 and whose value should always be a number. The class C1 also has a shared slot named S2.

There is a local slot named S1 in instances of C2. The default initial value of S1 is 5. The value of S1 will be of type (and integer number). There are also local slots named S2 and S3 in instances of C2. The class C2 has a method for C2-S3 for reading the value of slot S3; there is also a method for (setf C2-S3) that writes the value of S3.

Integrating Types and Classes

The Common Lisp Object System maps the space of classes into the Common Lisp type space. Every class that has a proper name has a corresponding type with the same name.

The proper name of every class is a valid type specifier. In addition, every class object is a valid type specifier. Thus the expression (typep object class) evaluates to true if the class of object is class itself or a subclass of class. The evaluation of the expression (subtypep class1 class2 ) returns the values t t if class1 is a subclass of class2 or if they are the same class; otherwise it returns the values nil t. If I is an instance of some class C named S and C is an instance of standard-class, the evaluation of the expression (type-of I) will return S if S is the proper name of C; if S is not the proper name of C, the expression (type-of I) will return C.

Because the names of classes and class objects are type specifiers, they may be used in the special form the and in type declarations.

Many but not all of the predefined Common Lisp type specifiers have a corresponding class with the same proper name as the type. These type specifiers are listed in Figure 1-1. For example, the type array has a corresponding class named array. No type specifier that is a list, such as (vector double-float 100), has a corresponding class. The form deftype does not create any classes.

Each class that corresponds to a predefined Common Lisp type specifier can be implemented in one of three ways, at the discretion of each implementation. It can be a standard class (of the kind defined by defclass), a structure class (defined by defstruct), or a built-in class (implemented in a special, non-extensible way).

A built-in class is one whose instances have restricted capabilities or special representations. Attempting to use defclass to define subclasses of a built-in class signals an error. Calling makeinstance to create an instance of a built-in class signals an error. Calling slot-value on an instance of a built-in class signals an error. Redefining a built-in class or using change-class to change the class of an instance to or from a built-in class signals an error. However, built-in classes can be used as parameter specializers in methods.

It is possible to determine whether a class is a built-in class by checking the metaclass. A standard class is an instance of standard-class, a built-in class is an instance of built-in-class, and a structure class is an instance of structure-class.

Each structure type created by defstruct without using the :type option has a corresponding class. This class is an instance of structure-class. The :include option of defstruct creates a direct subclass of the class that corresponds to the included structure.

The purpose of specifying that many of the standard Common Lisp type specifiers have a corresponding class is to enable users to write methods that discriminate on these types. Method selection requires that a class precedence list can be determined for each class.

The hierarchical relationships among the Common Lisp type specifiers are mirrored by relationships among the classes corresponding to those types. The existing type hierarchy is used for determining the class precedence list for each class that corresponds to a predefined Common Lisp type. In some cases, Common Lisp: The Language does not specify a local precedence order for two supertypes of a given type specifier. For example, null is a subtype of both symbol and list, but Common Lisp: The Language does not specify whether symbol is more specific or less specific than list. The Common Lisp Object System specification defines those relationships for all such classes.

The following figure lists the set of classes required by the Object System that correspond to predefined Common Lisp type specifiers. The superclasses of each such class are presented in order from most specific to most general, thereby defining the class precedence list for the class. The local precedence order for each class that corresponds to a Common Lisp type specifier can be derived from this table.

Predefined Common Lisp Type Class Precedence List for Corresponding Class
array  (array t)
bit-vector  (bit-vector vector array sequence t)
character  (character t)
complex  (complex number t)
cons  (cons list sequence t)
float  (float number t)
integer  (integer rational number t)
list  (list sequence t)
null  (null symbol list sequence t)
number  (number t)
ratio  (ratio rational number t)
rational  (rational number t)
sequence  (sequence t)
string  (string vector array sequence t)
symbol  (symbol t)
t  (t)
vector  (vector array sequence t)

Figure 1–1.

Individual implementations may be extended to define other type specifiers to have a corresponding class. Individual implementations can be extended to add other subclass relationships and to add other elements to the class precedence lists in the above table, as long as they do not violate the type relationships and disjointness requirements specified by Common Lisp: The Language. A standard class defined with no direct superclasses is guaranteed to be disjoint from all of the classes in the table, except for the class named t.

The following Common Lisp types will have corresponding classes when Common Lisp is modified to define them each as being disjoint from cons, symbol, array, number, and character:

• function

• hash-table

• package

• pathname

• random-state

• readtable

• stream

Determining the Class Precedence List

The defclass form for a class provides a total ordering on that class and its direct superclasses. This ordering is called the local precedence order. It is an ordered list of the class and its direct superclasses. The class precedence list for a class C is a total ordering on C and its superclasses that is consistent with the local precedence orders for each of C and its superclasses.

A class precedes its direct superclasses, and a direct superclass precedes all other direct superclasses specified to its right in the superclasses list of the defclass form. For every class C, define

R_{C} = {(C, C_{1}), (C_{1}, C_{2}), …, (C_{n−1}, C_{n})}

where C_{1}, …, C_{n} are the direct superclasses of C in the order in which they are mentioned in the defclass form. These ordered pairs generate the total ordering on the class C and its direct superclasses.

Let S_{C} be the set of C and its superclasses. Let R be

R = ⋃_{c∈S_{C}} R_{c}

The set R may or may not generate a partial ordering, depending on whether the R_{c}, c ∈ S_{C}, are consistent; it is assumed that they are consistent and that R generates a partial ordering. When the R_{c} are not consistent, it is said that R is inconsistent.

To compute the class precedence list for C, topologically sort the elements of S_{C} with respect to the partial ordering generated by R. When the topological sort must select a class from a set of two or more classes, none of which are preceded by other classes with respect to R, the class selected is chosen deterministically, as described below.

If R is inconsistent, an error is signaled.

Topological Sorting

Topological sorting proceeds by finding a class C in S_{C} such that no other class precedes that element according to the elements in R. The class C is placed first in the result. Remove C from S_{C}, and remove all pairs of the form (C, D), D ∈ S_{C}, from R. Repeat the process, adding classes with no predecessors to the end of the result. Stop when no element can be found that has no predecessor.

If S_{C} is not empty and the process has stopped, the set R is inconsistent. If every class in the finite set of classes is preceded by another, then R contains a loop. That is, there is a chain of classes C_{1}, …, C_{n} such that C_{i} precedes C_{i+1}, 1 ≤ i < n, and C_{n} precedes C_{1}.

Sometimes there are several classes from S_{C} with no predecessors. In this case select the one that has a direct subclass rightmost in the class precedence list computed so far. Because a direct superclass precedes all other direct superclasses to its right, there can be only one such candidate class. If there is no such candidate class, R does not generate a partial ordering—the R_{c}, c ∈ S_{C}, are inconsistent.

In more precise terms, let {N_{1}, …, N_{m}}, m ≥ 2, be the classes from S_{C} with no predecessors. Let (C_{1} … C_{n}), n ≥ 1, be the class precedence list constructed so far. C_{1} is the most specific class, and C_{n} is the least specific. Let 1 ≤ j ≤ n be the largest number such that there exists an i where 1 ≤ i ≤ m and N_{i} is a direct superclass of C_{j}; N_{i} is placed next.

The effect of this rule for selecting from a set of classes with no predecessors is that the classes in a simple superclass chain are adjacent in the class precedence list and that classes in each relatively separated subgraph are adjacent in the class precedence list. For example, let T_{1} and T_{2} be subgraphs whose only element in common is the class J. Suppose that no superclass of J appears in either T_{1} or T_{2}. Let C_{1} be the bottom of T_{1}; and let C_{2} be the bottom of T_{2}. Suppose C is a class whose direct superclasses are C_{1} and C_{2} in that order, then the class precedence list for C will start with C and will be followed by all classes in T_{1} except J. All the classes of T_{2} will be next. The class J and its superclasses will appear last.

Examples

This example determines a class precedence list for the class pie. The following classes are defined:

(defclass pie (apple cinnamon) ())
(defclass apple (fruit) ())
(defclass cinnamon (spice) ())
(defclass fruit (food) ())
(defclass spice (food) ())
(defclass food () ())
The set S = {pie, apple, cinnamon, fruit, spice, food, standard-object, t}. The set R =
{(pie, apple), (apple, cinnamon), (apple, fruit), (cinnamon, spice), (fruit, food),
(spice, food), (food, standard-object), (standard-object, t)}.

The class pie is not preceded by anything, so it comes first; the result so far is (pie). Remove

pie from S and pairs mentioning pie from R to get S = {apple, cinnamon, fruit, spice, food,
standard-object, t} and R = {(apple, cinnamon), (apple, fruit), (cinnamon, spice), (fruit,
food),
(spice, food), (food, standard-object), (standard-object, t)}.

The class apple is not preceded by anything, so it is next; the result is (pie apple). Removing

apple and the relevant pairs results in S = {cinnamon, fruit, spice, food, standard-object, t}
and R = {(cinnamon, spice), (fruit, food), (spice, food), (food, standard-object),
(standard-object, t)}.

The classes cinnamon and fruit are not preceded by anything, so the one with a direct subclass rightmost in the class precedence list computed so far goes next. The class apple is a direct subclass of fruit, and the class pie is a direct subclass of cinnamon. Because apple appears to the right of pie in the precedence list, fruit goes next, and the result so far is (pie apple fruit).

S = {cinnamon, spice, food, standard-object, t}; R = {(cinnamon, spice), (spice, food),
(food, standard-object), (standard-object, t)}.

The class cinnamon is next, giving the result so far as (pie apple fruit cinnamon). At this

point S = {spice, food, standard-object, t}; R = {(spice, food), (food, standard-object),
(standard-object, t)}.

The classes spice, food, standard-object, and t are added in that order, and the class precedence

list is (pie apple fruit cinnamon spice food standard-object t).

It is possible to write a set of class definitions that cannot be ordered. For example:

(defclass new-class (fruit apple) ())
(defclass apple (fruit) ())

The class fruit must precede apple because the local ordering of superclasses must be preserved. The class apple must precede fruit because a class always precedes its own superclasses. When this situation occurs, an error is signaled when the system tries to compute the class precedence list.

The following might appear to be a conflicting set of definitions:

(defclass pie (apple cinnamon) ())
(defclass pastry (cinnamon apple) ())
(defclass apple () ())
(defclass cinnamon () ())
The class precedence list for pie is (pie apple cinnamon standard-object t).
The class precedence list for pastry is (pastry cinnamon apple standard-object t).

It is not a problem for apple to precede cinnamon in the ordering of the superclasses of pie but not in the ordering for pastry. However, it is not possible to build a new class that has both pie and pastry as superclasses.

Generic Functions and Methods

A generic function is a function whose behavior depends on the classes or identities of the arguments supplied to it. The methods define the class-specific behavior and operations of the generic function. The following sections describe generic functions and methods.

Introduction to Generic Functions

A generic function object contains a set of methods, a lambda-list, a method combination type, and other information.

Like an ordinary Lisp function, a generic function takes arguments, performs a series of operations, and perhaps returns useful values. An ordinary function has a single body of code that is always executed when the function is called. A generic function has a set of bodies of code of which a subset is selected for execution. The selected bodies of code and the manner of their combination are determined by the classes or identities of one or more of the arguments to the generic function and by its method combination type.

Ordinary functions and generic functions are called with identical syntax.

Generic functions are true functions that can be passed as arguments and used as the first argument to funcall and apply.

In Common Lisp, a name can be given to an ordinary function in one of two ways: a global name can be given to a function using the defun construct; a local name can be given using the flet or labels special forms. A generic function can be given a global name using the defmethod or defgeneric construct. A generic function can be given a local name using the generic-flet, generic-labels, or with-added-methods special forms. The name of a generic function, like the name of an ordinary function, can be either a symbol or a two-element list whose first element is setf and whose second element is a symbol. This is true for both local and global names.

The generic-flet special form creates new local generic functions using the set of methods specified by the method definitions in the generic-flet form. The scoping of generic function names within a generic-flet form is the same as for flet.

The generic-labels special form creates a set of new mutually recursive local generic functions using the set of methods specified by the method definitions in the generic-labels form. The scoping of generic function names within a generic-labels form is the same as for labels.

The with-added-methods special form creates new local generic functions by adding the set of methods specified by the method definitions with a given name in the with-added-methods form to copies of the methods of the lexically visible generic function of the same name. If there is a lexically visible ordinary function of the same name as one of specified generic functions, that function becomes the method function of the default method for the new generic function of that name.

The generic-function macro creates an anonymous generic function with the set of methods specified by the method definitions in the generic-function form.

When a defgeneric form is evaluated, one of three actions is taken:

• If a generic function of the given name already exists, the existing generic function object is modified. Methods specified by the current defgeneric form are added, and any methods in the existing generic function that were defined by a previous defgeneric form are removed. Methods added by the current defgeneric form might replace methods defined by defmethod or defclass. No other methods in the generic function are affected or replaced.

• If the given name names a non-generic function, a macro, or a special form, an error is signaled.

• Otherwise a generic function is created with the methods specified by the method definitions in the defgeneric form.

Some forms specify the options of a generic function, such as the type of method combination it uses or its argument precedence order. These forms will be referred to as “forms that specify generic function options.” These forms are: defgeneric, generic-function, generic-flet, generic-labels, and with-added-methods.

Some forms define methods for a generic function. These forms will be referred to as “methoddefining forms.” These forms are: defgeneric, defmethod, generic-function, generic-flet, generic-labels, with-added-methods, and defclass. Note that all the method-defining forms except defclass and defmethod can specify generic function options and so are also forms that specify generic function options.

Introduction to Methods

A method object contains a method function, a sequence of parameter specializers that specify when the given method is applicable, a lambda-list, and a sequence of qualifiers that are used by the method combination facility to distinguish among methods.

A method object is not a function and cannot be invoked as a function. Various mechanisms in the Object System take a method object and invoke its method function, as is the case when a generic function is invoked. When this occurs it is said that the method is invoked or called.

A method-defining form contains the code that is to be run when the arguments to the generic function cause the method that it defines to be invoked. When a method-defining form is evaluated, a method object is created and one of four actions is taken:

• If a generic function of the given name already exists and if a method object already exists that agrees with the new one on parameter specializers and qualifiers, the new method object replaces the old one. For a definition of one method agreeing with another on parameter specializers and qualifiers, see the section “Agreement on Parameter Specializers and Qualifiers.”

• If a generic function of the given name already exists and if there is no method object that agrees with the new one on parameter specializers and qualifiers, the existing generic function object is modified to contain the new method object.

• If the given name names a non-generic function, a macro, or a special form, an error is signaled.

• Otherwise a generic function is created with the methods specified by the method-defining form.

If the lambda-list of a new method is not congruent with the lambda-list of the generic function, an error is signaled. If a method-defining form that cannot specify generic function options creates a new generic function, a lambda-list for that generic function is derived from the lambdalists of the methods in the method-defining form in such a way as to be congruent with them. For a discussion of congruence, see the section “Congruent Lambda-lists for All Methods of a Generic Function.”

Each method has a specialized lambda-list, which determines when that method can be applied. A specialized lambda-list is like an ordinary lambda-list except that a specialized parameter may occur instead of the name of a required parameter. A specialized parameter is a list (variable-name parameter-specializer-name), where parameter-specializer-name is one of the following:

• A name that names a class

• (eql form)

A parameter specializer name denotes a parameter specializer as follows:

• A name that names a class denotes that class.

• The list (eql form) denotes (eql object), where object is the result of evaluating form. The form form is evaluated in the lexical environment in which the method-defining form is evaluated. Note that form is evaluated only once, at the time the method is defined, not each time the generic function is called.

Parameter specializer names are used in macros intended as the user-level interface (defmethod), while parameter specializers are used in the functional interface.

Only required parameters may be specialized, and there must be a parameter specializer for each required parameter. For notational simplicity, if some required parameter in a specialized lambdalist in a method-defining form is simply a variable name, its parameter specializer defaults to the class named t.

Given a generic function and a set of arguments, an applicable method is a method for that generic function whose parameter specializers are satisfied by their corresponding arguments. The following definition specifies what it means for a method to be applicable and for an argument to satisfy a parameter specializer.

Let ⟨A_{1}, …, A_{n}⟩ be the required arguments to a generic function in order. Let ⟨P_{1}, …, P_{n}⟩ be the parameter specializers corresponding to the required parameters of the method M in order. The method M is applicable when each A_{i} satisfies P_{i}. If P_{i} is a class, and if A_{i} is an instance of a class C, then it is said that A_{i} satisfies P_{i} when C = P_{i} or when C is a subclass of P_{i}. If P_{i} is (eql object), then it is said that A_{i} satisfies P_{i} when the function eql applied to A_{i} and object is true.

Because a parameter specializer is a type specifier, the function typep can be used during method selection to determine whether an argument satisfies a parameter specializer. In general a parameter specializer cannot be a type specifier list, such as (vector single-float). The only parameter specializer that can be a list is (eql object). This requires that Common Lisp be modified to include the type specifier eql to be defined as if the following were evaluated:

(deftype eql (object) `(member ,object))

A method all of whose parameter specializers are the class named t is called a default method; it is always applicable but may be shadowed by a more specific method.

Methods can have qualifiers, which give the method combination procedure a way to distinguish among methods. A method that has one or more qualifiers is called a qualified method. A method with no qualifiers is called an unqualified method. A qualifier is any object other than a list, that is, any non-nil atom. The qualifiers defined by standard method combination and by the built-in method combination types are symbols.

In this specification, the terms primary method and auxiliary method are used to partition methods within a method combination type according to their intended use. In standard method combination, primary methods are unqualified methods and auxiliary methods are methods with a single qualifier that is one of :around, :before, or :after. When a method combination type is defined using the short form of define-method-combination, primary methods are methods qualified with the name of the type of method combination, and auxiliary methods have the qualifier :around. Thus the terms primary method and auxiliary method have only a relative definition within a given method combination type.

Agreement on Parameter Specializers and Qualifiers

Two methods are said to agree with each other on parameter specializers and qualifiers if the following conditions hold:

1. Both methods have the same number of required parameters. Suppose the parameter specializers of the two methods are P_{1,1} … P_{1,n} and P_{2,1} … P_{2,n}.

2. For each 1 ≤ i ≤ n, P_{1,i} agrees with P_{2,i}. The parameter specializer P_{1,i} agrees with P_{2,i} if P_{1,i} and P_{2,i} are the same class or if P_{1,i} = (eql object_{1}), P_{2,i} = (eql object_{2}), and (eql object_{1} object_{2}). Otherwise P_{1,i} and P_{2,i} do not agree.

3. The lists of qualifiers of both methods contain the same non-nil atoms in the same order. That is, the lists are equal.

Congruent Lambda-Lists for All Methods of a Generic Function

These rules define the congruence of a set of lambda-lists, including the lambda-list of each method for a given generic function and the lambda-list specified for the generic function itself, if given.

1. Each lambda-list must have the same number of required parameters.

2. Each lambda-list must have the same number of optional parameters. Each method can supply its own default for an optional parameter.

3. If any lambda-list mentions &rest or &key, each lambda-list must mention one or both of them.

4. If the generic function lambda-list mentions &key, each method must accept all of the keyword names mentioned after &key, either by accepting them explicitly, by specifying &allow-other-keys, or by specifying &rest but not &key. Each method can accept additional keyword arguments of its own. The checking of the validity of keyword names is done in the generic function, not in each method. A method is invoked as if the keyword argument pair whose keyword is :allow-other-keys and whose value is t were supplied, though no such argument pair will be passed.

5. The use of &allow-other-keys need not be consistent across lambda-lists. If &allow-other-keys is mentioned in the lambda-list of any applicable method or of the generic function, any keyword arguments may be mentioned in the call to the generic function.

6. The use of &aux need not be consistent across methods.

If a method-defining form that cannot specify generic function options creates a generic function, and if the lambda-list for the method mentions keyword arguments, the lambda-list of the generic function will mention &key (but no keyword arguments).

Keyword Arguments in Generic Functions and Methods

When a generic function or any of its methods mentions &key in a lambda-list, the specific set of keyword arguments accepted by the generic function varies according to the applicable methods. The set of keyword arguments accepted by the generic function for a particular call is the union of the keyword arguments accepted by all applicable methods and the keyword arguments mentioned after &key in the generic function definition, if any. A method that has &rest but not &key does not affect the set of acceptable keyword arguments. If the lambda-list of any applicable method or of the generic function definition contains &allow-other-keys, all keyword arguments are accepted by the generic function.

The lambda-list congruence rules require that each method accept all of the keyword arguments mentioned after &key in the generic function definition, by accepting them explicitly, by specifying &allow-other-keys, or by specifying &rest but not &key. Each method can accept additional keyword arguments of its own, in addition to the keyword arguments mentioned in the generic function definition.

If a generic function is passed a keyword argument that no applicable method accepts, an error is signaled.

For example, suppose there are two methods defined for width as follows:

(defmethod width ((c character-class) &key font) ...)
(defmethod width ((p picture-class) &key pixel-size) ...)

Assume that there are no other methods and no generic function definition for width. The evaluation of the following form will signal an error because the keyword argument :pixel-size is not accepted by the applicable method.

(width (make-instance `character-class :char #\Q)
       :font 'baskerville :pixel-size 10)

The evaluation of the following form will signal an error.

(width (make-instance `picture-class :glyph (glyph #\Q))
       :font 'baskerville :pixel-size 10)

The evaluation of the following form will not signal an error if the class named character-

picture-class is a subclass of both picture-class and character-class.
(width (make-instance `character-picture-class :char #\Q)
       :font 'baskerville :pixel-size 10)

Method Selection and Combination

When a generic function is called with particular arguments, it must determine the code to execute. This code is called the effective method for those arguments. The effective method is a combination of the applicable methods in the generic function. A combination of methods is a Lisp expression that contains calls to some or all of the methods. If a generic function is called and no methods apply, the generic function no-applicable-method is invoked.

When the effective method has been determined, it is invoked with the same arguments that were passed to the generic function. Whatever values it returns are returned as the values of the generic function.

Determining the Effective Method

The effective method is determined by the following three-step procedure:

1. Select the applicable methods.

2. Sort the applicable methods by precedence order, putting the most specific method first.

3. Apply method combination to the sorted list of applicable methods, producing the effective method.

Selecting the Applicable Methods

This step is described in the section “Introduction to Methods.”

Sorting the Applicable Methods by Precedence Order

To compare the precedence of two methods, their parameter specializers are examined in order. The default examination order is from left to right, but an alternative order may be specified by the :argument-precedence-order option to defgeneric or to any of the other forms that specify generic function options.

The corresponding parameter specializers from each method are compared. When a pair of parameter specializers are equal, the next pair are compared for equality. If all corresponding parameter specializers are equal, the two methods must have different qualifiers; in this case, either method can be selected to precede the other.

If some corresponding parameter specializers are not equal, the first pair of parameter specializers that are not equal determines the precedence. If both parameter specializers are classes, the more specific of the two methods is the method whose parameter specializer appears earlier in the class precedence list of the corresponding argument. Because of the way in which the set of applicable methods is chosen, the parameter specializers are guaranteed to be present in the class precedence list of the class of the argument.

If just one parameter specializer is (eql object), the method with that parameter specializer precedes the other method. If both parameter specializers are eql forms, the specializers must be the same (otherwise the two methods would not both have been applicable to this argument).

The resulting list of applicable methods has the most specific method first and the least specific method last.

Applying Method Combination to the Sorted List of Applicable Methods

In the simple case—if standard method combination is used and all applicable methods are primary methods—the effective method is the most specific method. That method can call the next most specific method by using the function call-next-method. The method that callnext-method will call is referred to as the next method. The predicate next-method-p tests whether a next method exists. If call-next-method is called and there is no next most specific method, the generic function no-next-method is invoked.

In general, the effective method is some combination of the applicable methods. It is defined by a Lisp form that contains calls to some or all of the applicable methods, returns the value or values that will be returned as the value or values of the generic function, and optionally makes some of the methods accessible by means of call-next-method. This Lisp form is the body of the effective method; it is augmented with an appropriate lambda-list to make it a function.

The role of each method in the effective method is determined by its method qualifiers and the specificity of the method. A qualifier serves to mark a method, and the meaning of a qualifier is determined by the way that these marks are used by this step of the procedure. If an applicable method has an unrecognized qualifier, this step signals an error and does not include that method in the effective method.

When standard method combination is used together with qualified methods, the effective method is produced as described in the section “Standard Method Combination.”

Another type of method combination can be specified by using the :method-combination option of defgeneric or of any of the other forms that specify generic function options. In this way this step of the procedure can be customized.

New types of method combination can be defined by using the define-method-combination macro.

The meta-object level also offers a mechanism for defining new types of method combination. The generic function compute-effective-method receives as arguments the generic function, the method combination object, and the sorted list of applicable methods. It returns the Lisp form that defines the effective method. A method for compute-effective-method can be defined directly by using defmethod or indirectly by using define-method-combination. A method combination object is an object that encapsulates the method combination type and options specified by the :method-combination option to forms that specify generic function options.

Implementation Note:

In the simplest implementation, the generic function would compute the effective method each time it was called. In practice, this will be too inefficient for some implementations. Instead, these implementations might employ a variety of optimizations of the three-step procedure. Some illustrative examples of such optimizations are the following:

• Use a hash table keyed by the class of the arguments to store the effective method.

• Compile the effective method and save the resulting compiled function in a table.

• Recognize the Lisp form as an instance of a pattern of control structure and substitute a closure that implements that structure.

• Examine the parameter specializers of all methods for the generic function and enumerate all possible effective methods. Combine the effective methods, together with code to select from among them, into a single function and compile that function. Call that function whenever the generic function is called.

Standard Method Combination

Standard method combination is supported by the class standard-generic-function. It is used if no other type of method combination is specified or if the built-in method combination type standard is specified.

Primary methods define the main action of the effective method, while auxiliary methods modify that action in one of three ways. A primary method has no method qualifiers.

An auxiliary method is a method whose method qualifier is :before, :after, or :around. Standard method combination allows no more than one qualifier per method; if a method definition specifies more than one qualifier per method, an error is signaled.

• A :before method has the keyword :before as its only qualifier. A :before method specifies code that is to be run before any primary methods.

• An :after method has the keyword :after as its only qualifier. An :after method specifies code that is to be run after primary methods.

• An :around method has the keyword :around as its only qualifier. An :around method specifies code that is to be run instead of other applicable methods but which is able to cause some of them to be run.

The semantics of standard method combination is as follows:

• If there are any :around methods, the most specific :around method is called. It supplies the value or values of the generic function.

• Inside the body of an :around method, call-next-method can be used to call the next method. When the next method returns, the :around method can execute more code, perhaps based on the returned value or values. The generic function no-next-method is invoked if call-next-method is used and there is no applicable method to call. The function next-method-p may be used to determine whether a next method exists.

• If an :around method invokes call-next-method, the next most specific :around method is called, if one is applicable. If there are no :around methods or if call-next-method is called by the least specific :around method, the other methods are called as follows:

– All the :before methods are called, in most-specific-first order. Their values are ignored. An error is signaled if call-next-method is used in a :before method.

– The most specific primary method is called. Inside the body of a primary method, callnext-method may be used to call the next most specific primary method. When that method returns, the previous primary method can execute more code, perhaps based on the returned value or values. The generic function no-next-method is invoked if call-next-method is used and there are no more applicable primary methods. The function next-method-p may be used to determine whether a next method exists. If call-next-method is not used, only the most specific primary method is called.

– All the :after methods are called in most-specific-last order. Their values are ignored. An error is signaled if call-next-method is used in an :after method.

• If no :around methods were invoked, the most specific primary method supplies the value or values returned by the generic function. The value or values returned by the invocation of call-next-method in the least specific :around method are those returned by the most specific primary method.

In standard method combination, if there is an applicable method but no applicable primary method, an error is signaled.

The :before methods are run in most-specific-first order while the :after methods are run in least-specific-first order. The design rationale for this difference can be illustrated with an example. Suppose class C_{1} modifies the behavior of its superclass, C_{2}, by adding :before and :after methods. Whether the behavior of the class C_{2} is defined directly by methods on C_{2} or is inherited from its superclasses does not affect the relative order of invocation of methods on instances of the class C_{1}. Class C_{1}’s :before method runs before all of class C_{2}’s methods. Class C_{1}’s :after method runs after all of class C_{2}’s methods.

By contrast, all :around methods run before any other methods run. Thus a less specific :around method runs before a more specific primary method.

If only primary methods are used and if call-next-method is not used, only the most specific method is invoked; that is, more specific methods shadow more general ones.

Declarative Method Combination

The macro define-method-combination defines new forms of method combination. It provides a mechanism for customizing the production of the effective method. The default procedure for producing an effective method is described in the section “Determining the Effective Method.” There are two forms of define-method-combination. The short form is a simple facility while the long form is more powerful and more verbose. The long form resembles defmacro in that the body is an expression that computes a Lisp form; it provides mechanisms for implementing arbitrary control structures within method combination and for arbitrary processing of method qualifiers. The syntax and use of both forms of define-method-combination are explained in Chapter 2.

Built-in Method Combination Types

The Common Lisp Object System provides a set of built-in method combination types. To specify that a generic function is to use one of these method combination types, the name of the method combination type is given as the argument to the :method-combination option to defgeneric or to the :method-combination option to any of the other forms that specify generic function options.

The names of the built-in method combination types are +, and, append, list, max, min, nconc, or, progn, and standard.

The semantics of the standard built-in method combination type was described in the section “Standard Method Combination.” The other built-in method combination types are called simple built-in method combination types.

The simple built-in method combination types act as though they were defined by the short form of define-method-combination. They recognize two roles for methods:

• An :around method has the keyword symbol :around as its sole qualifier. The meaning of :around methods is the same as in standard method combination. Use of the functions call-next-method and next-method-p is supported in :around methods.

• A primary method has the name of the method combination type as its sole qualifier. For example, the built-in method combination type and recognizes methods whose sole qualifier is and; these are primary methods. Use of the functions call-next-method and nextmethod-p is not supported in primary methods.

The semantics of the simple built-in method combination types is as follows:

• If there are any :around methods, the most specific :around method is called. It supplies the value or values of the generic function.

• Inside the body of an :around method, the function call-next-method can be used to call the next method. The generic function no-next-method is invoked if call-next-method is used and there is no applicable method to call. The function next-method-p may be used to determine whether a next method exists. When the next method returns, the :around method can execute more code, perhaps based on the returned value or values.

• If an :around method invokes call-next-method, the next most specific :around method is called, if one is applicable. If there are no :around methods or if call-next-method is called by the least specific :around method, a Lisp form derived from the name of the built-in method combination type and from the list of applicable primary methods is evaluated to produce the value of the generic function. Suppose the name of the method combination type is operator and the call to the generic function is of the form

(generic-function a_{1} … a_{n})

Let M_{1}, …, M_{k} be the applicable primary methods in order; then the derived Lisp form is

(operator ⟨M_{1} a_{1} … a_{n}⟩ … ⟨M_{k} a_{1} … a_{n}⟩)

If the expression ⟨M_{i} a_{1} … a_{n}⟩ is evaluated, the method M_{i} will be applied to the arguments a_{1} … a_{n}. For example, if operator is or, the expression ⟨M_{i} a_{1} … a_{n}⟩ is evaluated only if ⟨M_{j} a_{1} … a_{n}⟩, 1 ≤ j < i, returned nil.

The default order for the primary methods is :most-specific-first. However, the order can be reversed by supplying :most-specific-last as the second argument to the :methodcombination option.

The simple built-in method combination types require exactly one qualifier per method. An error is signaled if there are applicable methods with no qualifiers or with qualifiers that are not supported by the method combination type. An error is signaled if there are applicable :around methods and no applicable primary methods.

Meta-Objects

The implementation of the Object System manipulates classes, methods, and generic functions. The meta-object protocol specifies a set of generic functions defined by methods on classes; the behavior of those generic functions defines the behavior of the Object System. The instances of the classes on which those methods are defined are called meta-objects. Programming at the meta-object protocol level involves defining new classes of meta-objects along with methods specialized on these classes.

Metaclasses

The metaclass of an object is the class of its class. The metaclass determines the representation of instances of its instances and the forms of inheritance used by its instances for slot descriptions and method inheritance. The metaclass mechanism can be used to provide particular forms of optimization or to tailor the Common Lisp Object System for particular uses. The protocol for defining metaclasses is discussed in the chapter “The Common Lisp Object System Meta-Object Protocol.”

Standard Metaclasses

The Common Lisp Object System provides a number of predefined metaclasses. These include the classes standard-class, built-in-class, and structure-class:

• The class standard-class is the default class of classes defined by defclass.

• The class built-in-class is the class whose instances are classes that have special implementations with restricted capabilities. Any class that corresponds to a standard Common Lisp type specified in Common Lisp: The Language might be an instance of built-in-class. The predefined Common Lisp type specifiers that are required to have corresponding classes are listed in Figure 1-1. It is implementation dependent whether each of these classes is implemented as a built-in class.

• All classes defined by means of defstruct are instances of structure-class.

Standard Meta-objects

The Object System supplies a set of meta-objects, called standard meta-objects. These include the class standard-object and instances of the classes standard-method, standard-generic-function, and method-combination.

• The class standard-method is the default class of methods defined by the forms defmethod, defgeneric, generic-function, generic-flet, generic-labels, and withadded-methods.

• The class standard-generic-function is the default class of generic functions defined by the forms defmethod, defgeneric, generic-function, generic-flet, generic-labels, withadded-methods, and defclass.

• The class named standard-object is an instance of the class standard-class and is a superclass of every class that is an instance of standard-class except itself and structureclass.

• Every method combination object is an instance of a subclass of the class methodcombination.

Object Creation and Initialization

The generic function make-instance creates and returns a new instance of a class. The first argument is a class or the name of a class, and the remaining arguments form an initialization argument list.

The initialization of a new instance consists of several distinct steps, including the following: combining the explicitly supplied initialization arguments with default values for the unsupplied initialization arguments, checking the validity of the initialization arguments, allocating storage for the instance, filling slots with values, and executing user-supplied methods that perform additional initialization. Each step of make-instance is implemented by a generic function to provide a mechanism for customizing that step. In addition, make-instance is itself a generic function and thus also can be customized.

The Object System specifies system-supplied primary methods for each step and thus specifies a well-defined standard behavior for the entire initialization process. The standard behavior provides four simple mechanisms for controlling initialization:

• Declaring a symbol to be an initialization argument for a slot. An initialization argument is declared by using the :initarg slot option to defclass. This provides a mechanism for supplying a value for a slot in a call to make-instance.

• Supplying a default value form for an initialization argument. Default value forms for initialization arguments are defined by using the :default-initargs class option to defclass. If an initialization argument is not explicitly provided as an argument to make-instance, the default value form is evaluated in the lexical environment of the defclass form that defined it, and the resulting value is used as the value of the initialization argument.

• Supplying a default initial value form for a slot. A default initial value form for a slot is defined by using the :initform slot option to defclass. If no initialization argument associated with that slot is given as an argument to make-instance or is defaulted by :defaultinitargs, this default initial value form is evaluated in the lexical environment of the defclass form that defined it, and the resulting value is stored in the slot. The initform form for a local slot may be used when creating an instance, when updating an instance to conform to a redefined class, or when updating an instance to conform to the definition of a different class. The initform form for a shared slot may be used when defining or re-defining the class.

• Defining methods for initialize-instance and shared-initialize. The slot-filling behavior described above is implemented by a system-supplied primary method for initializeinstance which invokes shared-initialize. The generic function shared-initialize implements the parts of initialization shared by these four situations: when making an instance, when re-initializing an instance, when updating an instance to conform to a redefined class, and when updating an instance to conform to the definition of a different class. The systemsupplied primary method for shared-initialize directly implements the slot-filling behavior described above, and initialize-instance simply invokes shared-initialize.

Initialization Arguments

An initialization argument controls object creation and initialization. It is often convenient to use keyword symbols to name initialization arguments, but the name of an initialization argument can be any symbol, including nil. An initialization argument can be used in two ways: to fill a slot with a value or to provide an argument for an initialization method. A single initialization argument can be used for both purposes.

An initialization argument list is a list of alternating initialization argument names and values. Its structure is identical to a property list and also to the portion of an argument list processed for &key parameters. As in those lists, if an initialization argument name appears more than once in an initialization argument list, the leftmost occurrence supplies the value and the remaining occurrences are ignored. The arguments to make-instance (after the first argument) form an initialization argument list. Error-checking of initialization argument names is disabled if the keyword argument pair whose keyword is :allow-other-keys and whose value is non-nil appears in the initialization argument list.

An initialization argument can be associated with a slot. If the initialization argument has a value in the initialization argument list, the value is stored into the slot of the newly created object, overriding any :initform form associated with the slot. A single initialization argument can initialize more than one slot. An initialization argument that initializes a shared slot stores its value into the shared slot, replacing any previous value.

An initialization argument can be associated with a method. When an object is created and a particular initialization argument is supplied, the generic functions initialize-instance, sharedinitialize, and allocate-instance are called with that initialization argument’s name and value as a keyword argument pair. If a value for the initialization argument is not supplied in the initialization argument list, the method’s lambda-list supplies a default value.

Initialization arguments are used in four situations: when making an instance, when re-initializing an instance, when updating an instance to conform to a redefined class, and when updating an instance to conform to the definition of a different class.

Because initialization arguments are used to control the creation and initialization of an instance of some particular class, we say that an initialization argument is “an initialization argument for” that class.

Declaring the Validity of Initialization Arguments

Initialization arguments are checked for validity in each of the four situations that use them. An initialization argument may be valid in one situation and not another. For example, the systemsupplied primary method for make-instance defined for the class standard-class checks the validity of its initialization arguments and signals an error if an initialization argument is supplied that is not declared as valid in that situation.

There are two means for declaring initialization arguments valid.

• Initialization arguments that fill slots are declared as valid by the :initarg slot option to defclass. The :initarg slot option is inherited from superclasses. Thus the set of valid initialization arguments that fill slots for a class is the union of the initialization arguments that fill slots declared as valid by that class and its superclasses. Initialization arguments that fill slots are valid in all four contexts.

• Initialization arguments that supply arguments to methods are declared as valid by defining those methods. The keyword name of each keyword parameter specified in the method’s lambda-list becomes an initialization argument for all classes for which the method is applicable. Thus method inheritance controls the set of valid initialization arguments that supply arguments to methods. The generic functions for which method definitions serve to declare initialization arguments valid are as follows:

– Making an instance of a class: allocate-instance, initialize-instance, and sharedinitialize. Initialization arguments declared as valid by these methods are valid when making an instance of a class.

– Re-initializing an instance: reinitialize-instance and shared-initialize. Initialization arguments declared as valid by these methods are valid when re-initializing an instance.

– Updating an instance to conform to a redefined class: update-instance-for-redefined-class and shared-initialize. Initialization arguments declared as valid by these methods are valid when updating an instance to conform to a redefined class.

– Updating an instance to conform to the definition of a different class update-instance-for-different-class and shared-initialize. Initialization arguments declared as valid by these methods are valid when updating an instance to conform to conform to the definition of a different class.

The set of valid initialization arguments for a class is the set of valid initialization arguments that either fill slots or supply arguments to methods, along with the predefined initialization argument :allow-other-keys. The default value for :allow-other-keys is nil. The meaning of :allow-other-keys is the same as when it is passed to an ordinary function.

Defaulting of Initialization Arguments

A default value form can be supplied for an initialization argument by using the :defaultinitargs class option. If an initialization argument is declared valid by some particular class, its default value form might be specified by a different class. In this case :default-initargs is used to supply a default value for an inherited initialization argument.

The :default-initargs option is used only to provide default values for initialization arguments; it does not declare a symbol as a valid initialization argument name. Furthermore, the :defaultinitargs option is used only to provide default values for initialization arguments when making an instance.

The argument to the :default-initargs class option is a list of alternating initialization argument names and forms. Each form is the default value form for the corresponding initialization argument. The default value form of an initialization argument is used and evaluated only if that initialization argument does not appear in the arguments to make-instance and is not defaulted by a more specific class. The default value form is evaluated in the lexical environment of the defclass form that supplied it; the resulting value is used as the initialization argument’s value.

The initialization arguments supplied to make-instance are combined with defaulted initialization arguments to produce a defaulted initialization argument list. A defaulted initialization argument list is a list of alternating initialization argument names and values in which unsupplied initialization arguments are defaulted and in which the explicitly supplied initialization arguments appear earlier in the list than the defaulted initialization arguments. Defaulted initialization arguments are ordered according to the order in the class precedence list of the classes that supplied the default values.

There is a distinction between the purposes of the :default-initargs and the :initform options with respect to the initialization of slots. The :default-initargs class option provides a mechanism for the user to give a default value form for an initialization argument without knowing whether the initialization argument initializes a slot or is passed to a method. If that initialization argument is not explicitly supplied in a call to make-instance, the default value form is used, just as if it had been supplied in the call. In contrast, the :initform slot option provides a mechanism for the user to give a default initial value form for a slot. An :initform form is used to initialize a slot only if no initialization argument associated with that slot is given as an argument to make-instance or is defaulted by :default-initargs.

The order of evaluation of default value forms for initialization arguments and the order of evaluation of :initform forms are undefined. If the order of evaluation is important, initializeinstance or shared-initialize methods should be used instead.

Rules for Initialization Arguments

The :initarg slot option may be specified more than once for a given slot.

The following rules specify when initialization arguments may be multiply defined:

• A given initialization argument can be used to initialize more than one slot if the same initialization argument name appears in more than one :initarg slot option.

• A given initialization argument name can appear in the lambda-list of more than one initialization method.

• A given initialization argument name can appear both in an :initarg slot option and in the lambda-list of an initialization method.

If two or more initialization arguments that initialize the same slot are given in the arguments to make-instance, the leftmost of these initialization arguments in the initialization argument list supplies the value, even if the initialization arguments have different names.

If two or more different initialization arguments that initialize the same slot have default values and none is given explicitly in the arguments to make-instance, the initialization argument that appears in a :default-initargs class option in the most specific of the classes supplies the value. If a single :default-initargs class option specifies two or more initialization arguments that initialize the same slot and none is given explicitly in the arguments to make-instance, the leftmost in the :default-initargs class option supplies the value, and the values of the remaining default value forms are ignored.

Initialization arguments given explicitly in the arguments to make-instance appear to the left of defaulted initialization arguments. Suppose that the classes C_{1} and C_{2} supply the values of defaulted initialization arguments for different slots, and suppose that C_{1} is more specific than C_{2}; then the defaulted initialization argument whose value is supplied by C_{1} is to the left of the defaulted initialization argument whose value is supplied by C_{2} in the defaulted initialization argument list. If a single :default-initargs class option supplies the values of initialization arguments for two different slots, the initialization argument whose value is specified farther to the left in the default-initargs class option appears farther to the left in the defaulted initialization argument list.

If a slot has both an :initform form and an :initarg slot option, and the initialization argument is defaulted using :default-initargs or is supplied to make-instance, the captured :initform form is neither used nor evaluated.

The following is an example of the above rules:

  (defclass q () ((x :initarg a)))
  (defclass r (q) ((x :initarg b))
    (:default-initargs a 1 b 2))

Defaulted

Form  Initialization Argument List Contents of Slot

 (make-instance 'r)                       (a 1 b 2)                     1
 (make-instance 'r 'a 3)                  (a 3 b 2)                     3
 (make-instance 'r 'b 4)                  (b 4 a 1)                     4
 (make-instance 'r 'a 1 'a 2)           (a 1 a 2 b 2)                   1

Shared-Initialize

The generic function shared-initialize is used to fill the slots of an instance using initialization arguments and :initform forms when an instance is created, when an instance is re-initialized, when an instance is updated to conform to a redefined class, and when an instance is updated to conform to a different class. It uses standard method combination. It takes the following arguments: the instance to be initialized, a specification of a set of names of slots accessible in that instance, and any number of initialization arguments. The arguments after the first two must form an initialization argument list.

The second argument to shared-initialize may be one of the following:

• It can be list of slot names, which specifies the set of those slot names.

• It can be nil, which specifies the empty set of slot names.

• It can be the symbol t, which specifies the set of all of the slots.

There is a system-supplied primary method for shared-initialize whose first parameter specializer is the class standard-object. This method behaves as follows on each slot, whether shared or local:

• If an initialization argument in the initialization argument list specifies a value for that slot, that value is stored into the slot, even if a value has already been stored in the slot before the method is run. The affected slots are independent of which slots are indicated by the second argument to shared-initialize.

• Any slots indicated by the second argument that are still unbound at this point are initialized according to their :initform forms. For any such slot that has an :initform form, that form is evaluated in the lexical environment of its defining defclass form and the result is stored into the slot. For example, if a :before method stores a value in the slot, the :initform form will not be used to supply a value for the slot. If the second argument specifies a name that does not correspond to any slots accessible in the instance, the results are unspecified.

• The rules mentioned in the section “Rules for Initialization Arguments” are obeyed.

The generic function shared-initialize is called by the system-supplied primary methods for reinitialize-instance, update-instance-for-different-class, update-instance-for-redefined-class, and initialize-instance. Thus, methods can be written for shared-initialize to specify actions that should be taken in all of these contexts.

Initialize-Instance

The generic function initialize-instance is called by make-instance to initialize a newly created instance. It uses standard method combination. Methods for initialize-instance can be defined in order to perform any initialization that cannot be achieved with the simple slot-filling mechanisms.

During initialization, initialize-instance is invoked after the following actions have been taken:

• The defaulted initialization argument list has been computed by combining the supplied initialization argument list with any default initialization arguments for the class.

• The validity of the defaulted initialization argument list has been checked. If any of the initialization arguments has not been declared as valid, an error is signaled.

• A new instance whose slots are unbound has been created.

The generic function initialize-instance is called with the new instance and the defaulted initialization arguments. There is a system-supplied primary method for initialize-instance whose parameter specializer is the class standard-object. This method calls the generic function shared-initialize to fill in the slots according to the initialization arguments and the :initform forms for the slots; the generic function shared-initialize is called with the following arguments: the instance, t, and the defaulted initialization arguments.

Note that initialize-instance provides the defaulted initialization argument list in its call to shared-initialize, so the first step performed by the system-supplied primary method for shared-initialize takes into account both the initialization arguments provided in the call to make-instance and the defaulted initialization argument list.

Methods for initialize-instance can be defined to specify actions to be taken when an instance is initialized. If only :after methods for initialize-instance are defined, they will be run after the system-supplied primary method for initialization and therefore will not interfere with the default behavior of initialize-instance.

The Object System provides two functions that are useful in the bodies of initialize-instance methods. The function slot-boundp returns a boolean value that indicates whether a specified slot has a value; this provides a mechanism for writing :after methods for initializeinstance that initialize slots only if they have not already been initialized. The function slotmakunbound causes the slot to have no value.

Definitions of Make-Instance and Initialize-Instance

The generic function make-instance behaves as if it were defined as follows, except that certain optimizations are permitted:

(defmethod make-instance ((class standard-class) &rest initargs)
  (setq initargs (default-initargs class initargs))
  ...
  (let ((instance (apply #'allocate-instance class initargs)))
    (apply #'initialize-instance instance initargs)
    instance))
(defmethod make-instance ((class-name symbol) &rest initargs)
  (apply #'make-instance (find-class class-name) initargs))

The elided code in the definition of make-instance checks the supplied initialization arguments to determine whether an initialization argument was supplied that neither filled a slot nor supplied an argument to an applicable method. This check could be implemented using the generic functions class-prototype, compute-applicable-methods, function-keywords, and classslot-initargs. See Chapter 3 for a description of this initialization argument check.

The generic function initialize-instance behaves as if it were defined as follows, except that certain optimizations are permitted:

(defmethod initialize-instance ((instance standard-object) &rest initargs)
  (apply #'shared-initialize instance t initargs)))

These procedures can be customized at either the Programmer Interface level, the meta-object level, or both.

Customizing at the Programmer Interface level includes using the :initform, :initarg, and :default-initargs options to defclass, as well as defining methods for make-instance and initialize-instance. It is also possible to define methods for shared-initialize, which would be invoked by the generic functions reinitialize-instance, update-instance-for-redefined-class, update-instance-for-different-class, and initialize-instance. The meta-object level supports additional customization by allowing methods to be defined on make-instance, defaultinitargs, and allocate-instance. Chapters 2 and 3 document each of these generic functions and the system-supplied primary methods.

Implementations are permitted to make certain optimizations to initialize-instance and shared-initialize. The description of shared-initialize in Chapter 2 mentions the possible optimizations.

Because of optimization, the check for valid initialization arguments might not be implemented using the generic functions class-prototype, compute-applicable-methods, functionkeywords, and class-slot-initargs. In addition, methods for the generic function defaultinitargs, and the system-supplied primary methods for allocate-instance, initialize-instance, and shared-initialize might not be called on every call to make-instance or might not receive exactly the arguments that would be expected.

Redefining Classes

A class that is an instance of standard-class can be redefined if the new class will also be an instance of standard-class. Redefining a class modifies the existing class object to reflect the new class definition; it does not create a new class object for the class. Any method object created by a :reader, :writer, or :accessor option specified by the old defclass form is removed from the corresponding generic function. Methods specified by the new defclass form are added.

When the class C is redefined, changes are propagated to its instances and to instances of any of its subclasses. Updating such an instance occurs at an implementation-dependent time, but no later than the next time a slot of that instance is read or written. Updating an instance does not change its identity as defined by the eq function. The updating process may change the slots of that particular instance, but it does not create a new instance. Whether updating an instance consumes storage is implementation dependent.

Note that redefining a class may cause slots to be added or deleted. If a class is redefined in a way that changes the set of local slots accessible in instances, the instances will be updated. It is implementation dependent whether instances are updated if a class is redefined in a way that does not change the set of local slots accessible in instances.

The value of a slot that is specified as shared both in the old class and in the new class is retained. If such a shared slot was unbound in the old class, it will be unbound in the new class. Slots that were local in the old class and that are shared in the new class are initialized. Newly added shared slots are initialized.

Each newly added shared slot is set to the result of evaluating the captured :initform form for the slot that was specified in the defclass form for the new class. If there is no :initform form, the slot is unbound.

If a class is redefined in such a way that the set of local slots accessible in an instance of the class is changed, a two-step process of updating the instances of the class takes place. The process may be explicitly started by invoking the generic function make-instances-obsolete. This twostep process can happen in other circumstances in some implementations. For example, in some implementations this two-step process will be triggered if the order of slots in storage is changed.

The first step modifies the structure of the instance by adding new local slots and discarding local slots that are not defined in the new version of the class. The second step initializes the newly added local slots and performs any other user-defined actions. These two steps are further specified in the next two sections.

Modifying the Structure of Instances

The first step modifies the structure of instances of the redefined class to conform to its new class definition. Local slots specified by the new class definition that are not specified as either local or shared by the old class are added, and slots not specified as either local or shared by the new class definition that are specified as local by the old class are discarded. The names of these added and discarded slots are passed as arguments to update-instance-for-redefined-class as described in the next section.

The values of local slots specified by both the new and old classes are retained. If such a local slot was unbound, it remains unbound.

The value of a slot that is specified as shared in the old class and as local in the new class is retained. If such a shared slot was unbound, the local slot will be unbound.

Initializing Newly Added Local Slots

The second step initializes the newly added local slots and performs any other user-defined actions. This step is implemented by the generic function update-instance-for-redefined-class, which is called after completion of the first step of modifying the structure of the instance.

The generic function update-instance-for-redefined-class takes four required arguments: the instance being updated after it has undergone the first step, a list of the names of local slots that were added, a list of the names of local slots that were discarded, and a property list containing the slot names and values of slots that were discarded and had values. Included among the discarded slots are slots that were local in the old class and that are shared in the new class.

The generic function update-instance-for-redefined-class also takes any number of initialization arguments. When it is called by the system to update an instance whose class has been redefined, no initialization arguments are provided.

There is a system-supplied primary method for update-instance-for-redefined-class whose parameter specializer for its instance argument is the class standard-object. First this method checks the validity of initialization arguments and signals an error if an initialization argument is supplied that is not declared as valid. (See the section “Declaring the Validity of Initialization Arguments” for more information.) Then it calls the generic function shared-initialize with the following arguments: the instance, the list of names of the newly added slots, and the initialization arguments it received.

Customizing Class Redefinition

Methods for update-instance-for-redefined-class may be defined to specify actions to be taken when an instance is updated. If only :after methods for update-instance-for-redefined-class are defined, they will be run after the system-supplied primary method for initialization and therefore will not interfere with the default behavior of update-instance-for-redefined-class. Because no initialization arguments are passed to update-instance-for-redefined-class when it is called by the system, the initform forms for slots that are filled by :before methods for update-instance-for-redefined-class will not be evaluated by shared-initialize.

Methods for shared-initialize may be defined to customize class redefinition. See the section “Shared-Initialize” for more information.

Extensions

There are two allowed extensions to class redefinition:

• The Object System may be extended to permit the new class to be an instance of a metaclass other than the metaclass of the old class.

• The Object System may be extended to support an updating process when either the old or the new class is an instance of a class other than standard-class that is not a built-in class.

Changing the Class of an Instance

The function change-class can be used to change the class of an instance from its current class, C_{from}, to a different class, C_{to}; it changes the structure of the instance to conform to the definition of the class C_{to}.

Note that changing the class of an instance may cause slots to be added or deleted.

When change-class is invoked on an instance, a two-step updating process takes place. The first step modifies the structure of the instance by adding new local slots and discarding local slots that are not specified in the new version of the instance. The second step initializes the newly added local slots and performs any other user-defined actions. These two steps are further described in the two following sections.

Modifying the Structure of the Instance

In order to make the instance conform to the class C_{to}, local slots specified by the class C_{to} that are not specified by the class C_{from} are added, and local slots not specified by the class C_{to} that are specified by the class C_{from} are discarded.

The values of local slots specified by both the class C_{to} and the class C_{from} are retained. If such a local slot was unbound, it remains unbound.

The values of slots specified as shared in the class C_{from} and as local in the class C_{to} are retained.

This first step of the update does not affect the values of any shared slots.

Initializing Newly Added Local Slots

The second step of the update initializes the newly added slots and performs any other userdefined actions. This step is implemented by the generic function update-instance-for-different-class. The generic function update-instance-for-different-class is invoked by change-class after the first step of the update has been completed.

The generic function update-instance-for-different-class is invoked on two arguments computed by change-class. The first argument passed is a copy of the instance being updated and is an instance of the class C_{from}; this copy has dynamic extent within the generic function changeclass. The second argument is the instance as updated so far by change-class and is an instance of the class C_{to}.

The generic function update-instance-for-different-class also takes any number of initialization arguments. When it is called by change-class, no initialization arguments are provided.

There is a system-supplied primary method for update-instance-for-different-class that has two parameter specializers, each of which is the class standard-object. First this method checks the validity of initialization arguments and signals an error if an initialization argument is supplied that is not declared as valid. (See the section “Declaring the Validity of Initialization Arguments” for more information.) Then it calls the generic function shared-initialize with the following arguments: the instance, a list of names of the newly added slots, and the initialization arguments it received.

Customizing the Change of Class of an Instance

Methods for update-instance-for-different-class may be defined to specify actions to be taken when an instance is updated. If only :after methods for update-instance-for-different-class are defined, they will be run after the system-supplied primary method for initialization and will not interfere with the default behavior of update-instance-for-different-class. Because no initialization arguments are passed to update-instance-for-different-class when it is called by change-class, the initform forms for slots that are filled by :before methods for updateinstance-for-different-class will not be evaluated by shared-initialize.

Methods for shared-initialize may be defined to customize class redefinition. See the section “Shared-Initialize” for more information.

Reinitializing an Instance

The generic function reinitialize-instance may be used to change the values of slots according to initialization arguments.

The process of reinitialization changes the values of some slots and performs any user-defined actions. It does not modify the structure of an instance to add or delete slots, and it does not use any :initform forms to initialize slots.

The generic function reinitialize-instance may be called directly. It takes one required argument, the instance. It also takes any number of initialization arguments to be used by methods for reinitialize-instance or for shared-initialize. The arguments after the required instance must form an initialization argument list.

There is a system-supplied primary method for reinitialize-instance whose parameter specializer is the class standard-object. First this method checks the validity of initialization arguments and signals an error if an initialization argument is supplied that is not declared as valid. (See the section “Declaring the Validity of Initialization Arguments” for more information.) Then it calls the generic function shared-initialize with the following arguments: the instance, nil, and the initialization arguments it received.

Customizing Reinitialization

Methods for reinitialize-instance may be defined to specify actions to be taken when an instance is updated. If only :after methods for reinitialize-instance are defined, they will be run after the system-supplied primary method for initialization and therefore will not interfere with the default behavior of reinitialize-instance.

Methods for shared-initialize may be defined to customize class redefinition. See the section “Shared-Initialize” for more information.